OpenStudy (anonymous):

Monisha decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a joining fee of $100 and a monthly fee of $37. If x represents the number of months that Monisha is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time: 100 + 37x = C Monisha has allocated a maximum of $285 to spend on her gym membership. Which number line shows the possible number of months that Monisha can be a member of the gym?

OpenStudy (anonymous):

A number line shows numbers from 0 to 12 in increments of 1. The portion of the line from 1 to 5 is bold with shaded circles at 1 and at 5 A number line shows numbers from 0 to 12 in increments of 1. The portion of the line from 1 to 5 is bold with open circles at 1 and at 5 A number line shows numbers from 0 to 12 in increments of 1. The portion of the line from 0 to 5 is bold with shaded circles at 0 and at 5 A number line shows numbers from 0 to 12 in increments of 1. The portion of the line from 0 to 5 is bold with open circles at 0 and at 5

OpenStudy (anonymous):

@Mr.Question

OpenStudy (anonymous):

100 + 37x = C (where C = less then or equal to 285) 100 + 37x <=285 37x <= 285 - 100 37x <= 185 x <= 185/37 x <= 5 Now look at your number lines.....since the 5 is part of the solution, the 5 has to have a closed circle with shading to the left.all the way to 0. I believe that it is the third number line :)

OpenStudy (mathstudent55):

Since she has $285, let C = 285 and solve for x.